one number is 8 times a first number,a third number is 100 more than the first number. if the sum of the number is 440 find the numbers
Answer by oberobic(2304)

x = a number
y = another number
y = 8x
z = a third number
z = x+100
x +y +z = 440
.
substitute
.
x + 8x + x+100 = 440
10x + 100 = 440
10x = 340
x = 34
.
y = 8*34
y = 272
.
z = 34 + 100
z = 134
.
Check: Do they total 440?
.
34+272+134 = 440
Correct.
.
Answer:
The three numbers are 34, 272, and 134.
.
Done.